OpenStudy (anonymous):

Find the vertical asymptotes of the function below. x^2+8/ 9x-4x^2 x= (small value) x= (large value)

OpenStudy (anonymous):

9x-4x^2=x(9-4x)=0 x=0 x=9/4 at those two values we are dividing by 0 so we have a vertical asymptote.

OpenStudy (amistre64):

the top doesnt cancel any of the bottom; so all zeros in the bottom are vas